OpenStudy (anonymous):

which of the following quadratic equations have the solutions x=-2 and x=5 A.-1/13(x-2)(x+5)=0 B.(x-2)(x+5)=0 C.(x+2)(x-5)=0 D.-1/13(x+2)(x-5)=0

OpenStudy (anonymous):

@ARMYRANGERS can u help ?

OpenStudy (armyrangers):

Ok well what you need to do is find out witch one =0 when you plugin the X and Y

OpenStudy (anonymous):

ok

OpenStudy (armyrangers):

When you think you got it let me know and I will check it for you

OpenStudy (anonymous):

i think its b

OpenStudy (armyrangers):

(-2-2)(5+5)=0 (-4)(10)=0 -40=0 I would try again

OpenStudy (anonymous):

D

OpenStudy (armyrangers):

Yes good job

OpenStudy (anonymous):

what is the factorization of the polynomial graphed below . assume the coefficient of x^2 is 1 in the corresponding polynomial write each factor as a polynomial in descending order

OpenStudy (anonymous):